# CodePlan: Unlocking Reasoning Potential in Large Language Models by Scaling Code-form Planning

### v1

@@ Abstract
Despite the remarkable success of large language models (LLMs) on traditional natural language processing tasks, their planning ability remains a critical bottleneck in tackling complex multi-step reasoning tasks. Existing approaches mainly rely on prompting or task-specific fine-tuning, often suffering from weak robustness and cross-task generalization. To address the limitation, we introduce CodePlan, a scalable paradigm that empowers LLMs to generate and follow code-form plans—pseudocode that outlines high-level, structured reasoning processes. By leveraging the structured and versatile nature of code, CodePlan effectively captures the rich semantics and control flows inherent to sophisticated reasoning.
Importantly, CodePlan allows the automatic extraction of code-form plans from massive, wide-ranging text corpora without the need for curated, task-specific datasets. This enables it to scale up efficiently and improve reasoning capabilities across diverse scenarios. To train CodePlan, we construct a large-scale dataset of 2M examples that integrate code-form plans with standard prompt-response pairs from existing corpora. With minimal computation overhead during both training and inference, CodePlan achieves a 25.1% relative improvement compared with directly generating responses, averaged across 13 challenging multi-step reasoning benchmarks, spanning mathematical reasoning, symbolic reasoning, instruction-following, multi-hop QA, and decision-making tasks. Further analysis reveals CodePlan’s increasing performance gains on more complex reasoning tasks, as well as significant data efficiency thanks to its generalization ability [^1] Code and data are publicly available on https://github.com/thu-coai/CodePlan..
@@ 1 Introduction
With the rapid progress in pre-training and post-training (Brown, 2020; Ouyang et al., 2022; Chung et al., 2024), large language models (LLMs) have exhibited remarkable performance across a wide range of natural language processing (NLP) tasks. However, as LLMs are tasked with solving increasingly complex problems that require multi-step reasoning, such as mathematical problems (Cobbe et al., 2021; Hendrycks et al., 2021), multi-hop question-answering (Yang et al., 2018; Trivedi et al., 2022), and decision-making (Shridhar et al., 2020; Xie et al., 2024), their limited planning capability has become a critical bottleneck (Yang et al., 2023). As shown in Figure 1, LLMs often exhibit various failure modes in multi-step reasoning, such as repetitive steps, incoherent logic, focus drift, and early answering (Yao et al., 2022; Lanham et al., 2023). Effective planning can frame the reasoning procedure, thereby guiding LLMs through the intricate low-level steps to ultimately solve the tasks (Wang et al., 2023a).
@@
Delving deeper, LLMs’ planning deficiencies largely stem from the fact that massive pre-training text corpora often highly compress the underlying reasoning structures, thereby obscuring the latent, high-level planning signals that LLMs should learn (Zelikman et al., 2024). To remedy this challenge, current approaches mainly frame LLMs’ reasoning procedures through either advanced prompting techniques (Wei et al., 2022a; Yao et al., 2024) or task-specific fine-tuning (Zelikman et al., 2022; Guan et al., 2024). However, prompting approaches typically impose strict requirements on model inherent capabilities as well as carefully designed prompts (Anagnostidis & Bulian, 2024), while task-specific fine-tuning approaches limit the models’ ability to generalize to new domains. To surmount the above issues, we propose CodePlan, a novel, scalable paradigm that empowers LLMs to generate and follow code-form plans—pseudocode that serves as high-level, structured blueprints of the reasoning process. By leveraging the structured and versatile nature of code (Madaan et al., 2022; Wang et al., 2022), CodePlan effectively captures the rich semantics and control flows inherent to sophisticated reasoning. As exemplified in Figure 1, code naturally supports various common reasoning structures, including the hierarchical composition of multiple subtasks (function making and calling), iterative steps (for-loops), and conditional multi-branch steps (if-statements). Importantly, CodePlan allows the automatic extraction of code-form plans from massive, wide-ranging text corpora, bypassing the need for curated, task-specific datasets, as natural language text naturally compresses and embeds the underlying planning signals. This allows CodePlan to scale efficiently and improve reasoning capabilities across diverse reasoning tasks beyond specific ones like mathematical reasoning (Yu et al., 2023). Table 1 summarizes the advantages of CodePlan against prior studies on LLM reasoning.
@@
Plan-and-Solve Wang et al. (2023a) Free-Form Natural Language Text
@@
I. We introduce CodePlan, a novel, scalable paradigm that empowers LLMs to generate and follow code-form plans—pseudocode that outlines high-level, structured reasoning processes. This paradigm unlocks new frontiers for structured reasoning with LLMs, transcending the limitations imposed by the obscured implicit planning signals in natural language text.
II. CodePlan allows the efficient and cost-effective training data construction from massive, wide-ranging corpora, enabling promising data scalability. We exemplify this by curating a large-scale dataset comprised of 2M prompt-response pairs along with their corresponding code-form plans. This dataset also establishes a rich resource for future research on reasoning in LLMs.
III. We demonstrate the remarkable efficacy and generality of CodePlan across 13 challenging reasoning benchmarks on multiple backbone models, scaling from 7B to 13B. Further analysis reveals CodePlan’s growing superiority over baselines as problem complexity increases, as well as its superior data efficiency.
2 Method
We formally define the task requiring multi-step reasoning as follows: Given a natural language prompt $X$ that poses a problem, the goal is to generate a response $Y$ that provides a comprehensive solution through a sequence of logical reasoning steps. In this section, we elaborate on the methodology for solving the task by decomposing the generation process (§2.1) into two stages: planning (§2.2) and surface realization (§2.3).
@@ 2.1 Formalization
$\mathcal{L}=-\log~{}p(Y|X),$ (1)
To imbue LLMs with structured and systematic reasoning capabilities, we propose decomposing the generation process into two stages: planning and surface realization (Reiter & Dale, 2000). The planning stage aims to generate a structured plan $Z$ that outlines the key steps and control flow for solving the problem $X$, while the surface realization stage then translates this plan into the final natural language response $Y$, fleshing out the low-level reasoning details. The optimization objective $\mathcal{L}$ is then modeled as follows:
$\mathcal{L}=-\log p(Y|X)=-\log\mathbb{E}_{p(Z|X)}p(Y|X,Z),$ (2)
where $p(Z|X)$ and $p(Y|X,Z)$ refer to the models for planning and surface realization, respectively. Nevertheless, marginalizing over the latent variable $Z$ is generally intractable, as the search space could be vast. To circumvent the challenge, we minimize a variational upper bound of the loss with a posterior $q(Z|X,Y)$. Hence, the objective can be formulated as
$\mathcal{L}$ $=-\log\mathbb{E}_{p(Z|X)}p(Y|X,Z)\leqslant-\mathbb{E}_{q(Z|X,Y)}\log p(Y|X,Z)+\text{KL}\left(q(Z|X,Y)||p(Z|X)\right).$ (3)
@@
(2) Implicit Plans. Alternatively, plans can be modeled as learnable latent vectors that implicitly encode essential factors such as expert knowledge, styles, and other nuances shaping the reasoning process (Kingma, 2013). In this context, $q(Z|X,Y)$ can be defined as continuous Gaussian distribution (Kingma, 2013), discrete multinomial distribution over a learnable codebook (Van Den Oord et al., 2017; Cornille et al., 2024) or a pre-defined vocabulary (Zelikman et al., 2024). Compared to explicit representations, latent vectors have the potential to capture more subtle planning aspects when forming $Y$ given $X$, provided that the latent codes are learned from data. However, this approach may suffer from posterior collapse (Bowman et al., 2016) and computational inefficiency.
In this work, we adopt the simple and explainable “Explicit Plans” approach by setting $q(Z|X,Y)=\delta(Z=Z^{*})$, where $Z^{*}$ denotes the plan provided by a “gold annotator” due to its simplicity and scalability. We reserve the exploration of the other approach for future work. The “Explicit Plans” setting then yields the following variational upper bound for Eq. 3:
@@
For simplicity, we unify the planning model $p(Z^{*}|X)$ and the surface realization model $p(Y|X,Z^{*})$ into a single LM architecture, parameterized by shared parameters $\theta$. Importantly, this formulation implies that an optimal plan $Z^{*}$ should achieve a delicate balance: it must be sufficiently informative to guide the reasoning process effectively, yet concise enough to be easily and efficiently generated. This balance is critical for minimizing the combined difficulty of plan generation and subsequent realization, thereby optimizing overall model performance.
@@ 2.2 Planning
Given an input prompt $X$, $Z^{*}$ should capture the rich semantics and control flows inherent in the multi-step reasoning path $Y$. This motivates us to use programming languages—which are Turing complete (Boyer & Moore, 1983)—as a general representation of $Z^{*}$, thereby framing planning generation as code generation.
As exemplified in Figure 1, this code-form representation is versatile across diverse scenarios with several compelling advantages: (1) It seamlessly incorporates conditional branching (i.e., if-else statements) to handle diverse scenarios, where the reasoning step should dynamically adapt to intermediate results or contexts. (2) It integrates iterative loops (i.e., for blocks) to handle sequential data or perform repeated operations. (3) It defines and composes modular tools (i.e., Python functions), thereby enhancing LLMs’ abilities to craft and use tools (Cai et al., 2023). This also potentially endows our approach with the ability to build agents that can interact with external environments through specific APIs (Hausknecht et al., 2020), which are left for future work. (4) From a high level perspective, it naturally supports a hierarchical reasoning structure. It defines variables and attributes upfront, addresses sub-tasks via specific functions, and orchestrates the main procedure using rigorous formal logic. This hierarchical decomposition fosters the breakdown of complex reasoning problems into modular sub-components, facilitating the effective transfer of planning knowledge and enhancing the model’s systematic reasoning abilities (Yang et al., 2023).
Given a dataset of $(X,Y)$ pairs, we obtain the annotation of $Z^{*}$ by prompting an LLM that has been extensively pre-trained on code data, as detailed in §3.1. Specifically, we instruct the model to generate a short Python script that outlines the high-level reasoning structures for solving the problem $X$ and arriving at the response $Y$. Notably, we do not mandate the generated plans to be self-contained or executable, which is motivated by three key considerations: (1) leveraging code as a general and scalable representation for encoding reasoning logic is sufficient, and executing the generated code is not strictly necessary for our objective of enhancing LLMs’ reasoning capabilities; (2) generating executable codes is much less token efficient than generating pseudocode; and (3) our pilot study finds that it is still challenging for existing LLMs to generate fully executable code plans for various tasks, as this requires a comprehensive understanding of various libraries, APIs, and domain-specific knowledge. By circumventing the execution step, we can focus on the core challenge of learning to generate structured plans that capture the underlying reasoning logic, without being hindered by the additional complexities. After obtaining the dataset of $(X,Z^{*},Y)$ triples, we optimize the model for plan generation by minimizing the second term of Eq. 4 (i.e., $-\log p(Z^{*}|X)$).
@@ 2.3 Surface Realization
As illustrated in Figure 1, this enables the controllable generation of $Y$ that adheres to the human-readable plan $Z$. For instance, when realizing if-else blocks, the model is explicitly conditioned on the multi-branch conditions specified in the code, ensuring adherence to the intended logic. Similarly, when expanding for loops, the model is guided to systematically process each element, following the encoded iteration logic, which is critical for tasks involving iterative reasoning or sequential decision-making (Shridhar et al., 2020; Xie et al., 2024). This tight coupling between the code-form plan and the final reasoning path endows the model with the ability to produce coherent, logically sound solutions that faithfully reflect the intricate reasoning structure.
@@ 3.1 Training Data Curation
To facilitate the effective learning of planning and surface realization, we curate a large-scale training dataset comprised of numerous examples in the form $(X,Z^{*},Y)$. We leverage the WebInstruct dataset (Yue et al., 2024) as our starting point, which is automatically constructed from raw web data and contains a diverse array of natural language instructions (i.e., $X$) paired with high-quality responses (i.e., $Y$). We conduct zero-shot prompting based on Llama-3-8B-Instruct (Dubey et al., 2024) to automatically synthesize the corresponding gold plan $Z^{*}$ for each $(X,Y)$ pair. Table 2 shows the prompt. We use nucleus sampling (Holtzman et al., 2020) ($p=0.9$) with a temperature of 0.7. This automated data curation approach enables us to construct a large-scale dataset, and is readily extensible to other existing data sources such as Li et al. (2024) and (Cheng et al., 2024). More details are presented in Appendix A.1.
@@ 3.2 Baselines
We evaluate the following baselines: (1) Plan-and-Solve (PS) Prompting: It prompts the LLM to first devise a natural language plan to decompose the entire task into smaller subtasks, and then generate the response according to the plan (Wang et al., 2023b). (2) Vanilla Training: It first trains the LLM on the same prompt-response pairs used in CodePlan while eliminating the code-form plans, and then prompts the model to directly generate the response for evaluation.
@@ 3.3 Experimental Setup
We select Mistral-7B (Jiang et al., 2023) and Llama-2-7B/13B (Touvron et al., 2023) as our backbone models [^2] As there is no official model of $\sim$13B parameters in the Llama-3 series, we chose to conduct our experiments on the Llama-2 series models.. For Mistral-7B, we use a learning rate of 5e-6 and a global batch size of 512. For the Llama series, we use a learning rate of 1e-5 and a global batch of 256. We set the maximum training epochs to 2. To improve the training efficiency, we adopt mixed-precision training (Micikevicius et al., 2017), gradient checkpointing (Chen et al., 2016), FlashAttention2 (Dao, 2024), and ZeRO implemented in Deepspeed (Rasley et al., 2020). During inference, models are instructed to generate CoT-style responses for all tasks, elucidating their step-by-step rationale. Unless stated otherwise, we conduct evaluations under the few-shot setting (from 2-shot to 4-shot) without fine-tuning on the training sets of evaluation benchmarks. The above settings are also applied on baselines for fair comparisons. All our experiments are carried out on 32 A100 GPUs.

### v2

# Unlocking Reasoning Potential in Large Language Models by Scaling Code-form Planning

@@ Abstract
Despite the remarkable success of large language models (LLMs) on traditional natural language processing tasks, their planning ability remains a critical bottleneck in tackling complex multi-step reasoning tasks. Existing approaches mainly rely on prompting or task-specific fine-tuning, often suffering from poor robustness and cross-task generalization. To address the limitation, we introduce CodePlan, a scalable framework that empowers LLMs to generate and follow code-form plans—pseudocode that outlines high-level, structured reasoning processes. By leveraging the structured and versatile nature of code, CodePlan effectively captures the rich semantics and control flows inherent to sophisticated reasoning tasks. Importantly, CodePlan allows automatic extraction of code-form plans from massive, wide-ranging text corpora without the need for curated, task-specific datasets. This enables it to scale up efficiently and improve LLM’s reasoning capabilities across diverse scenarios. To train CodePlan, we construct a large-scale dataset of 2M examples that integrate code-form plans with standard prompt-response pairs from existing corpora. With minimal computation overhead during both training and inference, CodePlan achieves a 25.1% relative improvement compared with directly generating responses, averaged across 13 challenging multi-step reasoning benchmarks, spanning mathematical reasoning, symbolic reasoning, instruction-following, multi-hop QA, and decision-making tasks. Further analysis reveals CodePlan’s increasing performance gains on more complex reasoning tasks, as well as significant data efficiency thanks to its generalization ability.
@@ 1 Introduction
With the rapid progress in pre-training and post-training (Brown, 2020; Chung et al., 2024), large language models (LLMs) have exhibited remarkable performance across a wide range of natural language processing (NLP) tasks. However, as LLMs are tasked with solving increasingly complex problems that require multi-step reasoning, such as mathematical problems (Cobbe et al., 2021; Hendrycks et al., 2021), multi-hop question-answering (Trivedi et al., 2022), and complex decision-making (Shridhar et al., 2020; Xie et al., 2024), their limited planning capability has become a critical bottleneck (Yang et al., 2023). As illustrated in Figure 1, LLMs often exhibit various failure modes in multi-step reasoning, such as repetitive steps, incoherent logic, focus drift, and early answering (Yao et al., 2022; Lanham et al., 2023). Effective planning, i.e., generating a high-level abstraction in advance (Yang, 2012; Saitta et al., 2013; Russell & Norvig, 2016), can frame the subsequent reasoning procedure, thereby guiding LLMs through the intricate low-level steps to ultimately solve the tasks (Wang et al., 2023).
@@
Delving deeper, LLMs’ planning deficiencies largely stem from the fact that massive pre-training text corpora often do not explicitly exhait the underlying reasoning structures, thereby obscuring the latent, high-level planning signals that LLMs should learn (Zelikman et al., 2024). To remedy this challenge, current approaches mainly frame LLMs’ reasoning procedures through either advanced prompting techniques (Wei et al., 2022a; Yao et al., 2024) or task-specific fine-tuning (Zelikman et al., 2022; Guan et al., 2024). However, prompting approaches typically impose strict requirements on models’ inherent capabilities as well as carefully designed prompts (Anagnostidis & Bulian, 2024), while task-specific fine-tuning approaches limit the models’ ability to generalize to new domains.
To surmount the aforementioned issues, we propose CodePlan, a novel, scalable framework that empowers LLMs to generate and follow code-form plans—pseudocode that serves as high-level, structured blueprints of the reasoning process. By leveraging the structured and versatile nature of code (Madaan et al., 2022), CodePlan effectively captures the rich semantics and control flows inherent to sophisticated reasoning. As exemplified in Figure 1, code naturally supports various common reasoning structures, including the hierarchical composition of multiple subtasks (function making and calling), iterative steps (for-loops), and conditional multi-branch steps (if-statements). Importantly, CodePlan allows the automatic extraction of code-form plans from massive, wide-ranging text corpora that naturally embed the planning signals, bypassing the need for curated, task-specific datasets. This enables CodePlan to scale efficiently and improve reasoning capabilities across diverse tasks beyond specific ones like mathematical reasoning (Yu et al., 2023). Table 1 summarizes the advantages of CodePlan against prior work.
@@
Plan-and-Solve Wang et al. (2023) Free-Form Natural Language Text
@@
I. We introduce CodePlan, a novel, scalable framework that empowers LLMs to generate and follow code-form plans—pseudocode that outlines high-level, structured reasoning processes. This framework unlocks new frontiers for structured reasoning with LLMs, transcending the limitations imposed by the obscured implicit planning signals in natural language text.
II. CodePlan allows efficient and cost-effective training data construction from massive, wide-ranging corpora, enabling promising data scalability. We exemplify this by curating a large-scale dataset comprised of 2M prompt-response pairs along with their corresponding code-form plans. This dataset also establishes a rich resource for future research on reasoning in LLMs.
III. We demonstrate CodePlan’s remarkable efficacy and generality across 13 challenging reasoning benchmarks on multiple backbone models, scaling from 7B to 13B. Further analysis reveals its growing advantage over baselines as problem complexity increases, and its strong data efficiency.
2 Methodology
We formally define the multi-step reasoning task as follows: Given a prompt $X$ that poses a problem, the goal is to generate a response $Y$ that requires a comprehensive solution through a sequence of logical reasoning steps. In this section, we elaborate on the method for solving the task by decomposing the generation process (§2.1) into two stages: planning (§2.2) and surface realization (§2.3).
@@ 2.1 Formalization
$\mathcal{L}=-\log~{}p(Y|X).$ (1)
To imbue LLMs with structured and systematic reasoning capabilities, we propose decomposing the generation process into two stages: planning and surface realization (Reiter & Dale, 2000). The planning stage aims to generate a plan $Z$ that outlines the control flow for solving the problem $X$, while the surface realization stage then translates this plan into the final natural language response $Y$, fleshing out the low-level reasoning details. The optimization objective $\mathcal{L}$ is then modeled as follows:
$\mathcal{L}=-\log p(Y|X)=-\log\mathbb{E}_{p(Z|X)}p(Y|X,Z).$ (2)
where $p(Z|X)$ and $p(Y|X,Z)$ refer to the planning model and surface realization model, respectively. Nevertheless, marginalizing over the latent variable $Z$ is generally intractable, as the search space could be vast. To circumvent the challenge, we minimize a variational upper bound of the loss with a posterior $q(Z|X,Y)$ (Kingma, 2013). Hence, the objective can be formulated as:
$\mathcal{L}$ $=-\log\mathbb{E}_{p(Z|X)}p(Y|X,Z)\leqslant-\mathbb{E}_{q(Z|X,Y)}\log p(Y|X,Z)+\text{KL}\left(q(Z|X,Y)||p(Z|X)\right).).$ (3)
@@
(2) Implicit Plans. Alternatively, plans can be modeled as learnable latent vectors that implicitly encode expert knowledge, styles, or other nuances shaping the reasoning process (Kingma, 2013). In this context, $q(Z|X,Y)$ can be defined as continuous Gaussian distribution (Kingma, 2013), discrete multinomial distribution over a learnable codebook (Van Den Oord et al., 2017; Cornille et al., 2024) or a pre-defined vocabulary (Zelikman et al., 2024). Compared to explicit representations, latent vectors might capture more subtle planning aspects. However, this approach often introduces computation overhead and may suffer from posterior collapse (Bowman et al., 2016).
In this work, we adopt the simple and explainable “Explicit Plans” approach by setting $q(Z|X,Y)=\delta(Z=Z^{*})$, where $Z^{*}$ denotes the plan provided by a “gold annotator.” We reserve the exploration of the implicit approach for future work. This setting then yields the following upper bound for Eq. 3:
@@
For simplicity, we unify the planning model $p(Z^{*}|X)$ and the surface realization model $p(Y|X,Z^{*})$ into a single LM architecture with shared parameters $\theta$. Crucially, this formulation implies that an optimal plan $Z^{*}$ should achieve a delicate balance between informativeness for effective reasoning guidance and conciseness for efficient generation. This balance minimizes the combined difficulty of plan generation and subsequent realization, thereby optimizing overall model performance.
@@ 2.2 Planning
Given a prompt $X$, $Z^{*}$ should capture the rich semantics and control flows inherent in the reasoning path $Y$. This motivates us to use programming languages—which are Turing complete (Boyer & Moore, 1983)—as a general representation of $Z^{*}$, thereby framing planning as code generation.
As shown in Figure 1, this code-form representation is versatile across diverse scenarios with several compelling advantages: (1) It seamlessly incorporates conditional branching (i.e., if-statements) to dynamically adapt reasoning steps to intermediate results or contexts. (2) It integrates iterative loops (i.e., for-loops) to handle sequential data or perform repeated operations. (3) It defines and composes modular tools (i.e., Python functions), enhancing LLMs’ abilities to craft and use tools (Cai et al., 2024). This also potentially endows our approach with the ability to build agents that can interact with external environments through specific APIs (Hausknecht et al., 2020), which is left for future work. (4) From a high-level perspective, it naturally supports a hierarchical reasoning structure, by defining variables and attributes upfront, addressing sub-tasks via specific functions, and orchestrating the main procedure using rigorous formal logic. This fosters the breakdown of complex reasoning problems into modular sub-components, facilitating the effective transfer of planning knowledge and enhancing the model’s systematic reasoning abilities (Yang et al., 2023).
Given a dataset of $(X,Y)$ pairs, we obtain the annotation of $Z^{*}$ by prompting an LLM that has been extensively pre-trained on code data, as detailed in §3.1. Specifically, we instruct the model to generate a Python-style pseudocode that outlines the reasoning structure for solving the problem $X$ and arriving at the response $Y$. We do not mandate the generated plans to be executable in light of three considerations: (1) encoding reasoning logic using code is sufficient for generality and scalability, and execution might be unnecessary; (2) generating pseudocode is more token-efficient than executable code; and (3) our pilot study finds that existing LLMs still struggle to generate fully executable code plans for various tasks, as this requires a comprehensive understanding of various libraries, APIs, and domain-specific knowledge. By circumventing the execution step, we can focus on the core challenge of generating structured plans that capture the underlying reasoning logic, without additional complexities. After obtaining the dataset of $(X,Z^{*},Y)$ triples, we optimize the model for plan generation by minimizing the second term of Eq. 4 (i.e., $-\log p(Z^{*}|X)$).
@@ 2.3 Surface Realization
As illustrated in Figure 1, this enables the controllable generation of $Y$ that adheres to the human-readable plan $Z$. For instance, when realizing if-statements, the model is explicitly conditioned on the multi-branch conditions specified in the code, ensuring adherence to the intended logic. Similarly, when expanding for-loops, the model is guided to systematically process each element, following the encoded iteration logic, which is critical for tasks involving iterative reasoning or sequential decision-making (Shridhar et al., 2020; Xie et al., 2024). This tight coupling between the code-form plan and the final reasoning path enables the model to produce coherent, logically sound solutions that faithfully reflect the intricate reasoning structure.
@@ 3.1 Training Data Curation
To facilitate effective learning of planning and surface realization, we curate a large-scale training dataset of examples in the form $(X,Z^{*},Y)$. We start from WebInstruct (Yue et al., 2024) that is automatically constructed from raw web data and contains diverse prompt-response pairs, and prompt Llama-3-8B-Instruct (Dubey et al., 2024) to synthesize $Z^{*}$. Table 2 shows the prompt. We use nucleus sampling (Holtzman et al., 2020) ($p=0.9$) with a temperature of 0.7. This approach enables efficient construction of large-scale datasets, and is readily extensible to other existing corpora such as Li et al. (2024) and Cheng et al. (2024b). More details are presented in Appendix A.1.
@@ 3.2 Baselines
We evaluate the following baselines: (1) Plan-and-Solve (PS) Prompting: It prompts the LLM to first devise a natural language plan to decompose the entire task into smaller subtasks, and then generate the response following the plan (Wang et al., 2023). (2) Quiet-STaR: It automatically learns implicit plans for generating each token from WebInstruct (Zelikman et al., 2024). (3) Vanilla Training: It first trains the LLM on WebInstruct and then prompts it to directly generate the response.
@@ 3.3 Experimental Setup
We select Mistral-7B (Jiang et al., 2023) and Llama-2-7B/13B (Touvron et al., 2023) as our backbone models [^1] As there is no official $\sim$13B model in the Llama-3 series, we conducted experiments on Llama-2 models.. For Mistral/Llama, we use a learning rate of 5e-6/1e-5 and a global batch size of 512/256, respectively. We set the maximum training epochs to 2. To improve the training efficiency, we adopt mixed-precision training (Micikevicius et al., 2017), gradient checkpointing (Chen et al., 2016), FlashAttention2 (Dao, 2024), and ZeRO implemented in Deepspeed (Rasley et al., 2020). During inference, models are instructed to generate CoT-style responses for all tasks. Unless stated otherwise, we conduct evaluations under the few-shot setting (from 2-shot to 4-shot) without fine-tuning on evaluation benchmarks. The above settings are also applied on baselines for fair comparisons.